OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u=<6,9> v=3/2,-1>

OpenStudy (anonymous):

take u dot v =??

OpenStudy (anonymous):

I got 5/2

OpenStudy (anonymous):

\[u\bullet v=6*\dfrac{3}{2}+9*(-1)=9-9 =0\] How can you get 5/2? and u dot v=0--> u\(\perp\) v